What is the value of p in the equation
(p-5)/12 = (p-3)/7?
Which answer is it?
a. 1/5
b. 1/4
c. 2/3
d. 1/2

Answer by Alan3354(69443)

What is the value of p in the equation
(p-5)/12 = (p-3)/7?
------------------
Cross multiply
7(p-5) = 12(p-3)